OpenStudy (anonymous):

okay...the table lists data regarding the avg saleries of athletes in the years 1991 ($272,000) and 2001 ( 1,450,000). What is the linear function that fits the data S(x) =? What is the predicted avg salary for 2005? For 2010?

OpenStudy (saifoo.khan):

what if i say thats easy?

OpenStudy (anonymous):

I would tell you that you are nuts. LOL

OpenStudy (saifoo.khan):

HAHA.

OpenStudy (anonymous):

You are funny...that is not easy.

OpenStudy (anonymous):

oh crap, let me get my glasses on...

OpenStudy (saifoo.khan):

HAHA1.

OpenStudy (saifoo.khan):

should i solve futher?

OpenStudy (saifoo.khan):

further*

OpenStudy (dumbcow):

use slope formula \[m = \frac{y2-y1}{x2-x1} = \frac{1450000 - 272000}{2001-1991} = \frac{1178000}{10} = 117800\]

OpenStudy (anonymous):

yea, what is the predicted salary for 2005 and for 2010?

OpenStudy (anonymous):

wow, dumbcow...you are intense.

OpenStudy (anonymous):

So 117, 800 is the salary for what year?

OpenStudy (dumbcow):

that is the slope, it increases by that amout each year S(x) = 272,000 + 117,800x where x is years after 1991 2005: x = 14 2010: x =19

OpenStudy (anonymous):

Okay, that would be for 14 years right?

OpenStudy (anonymous):

117,800 would be for the 14 years????

OpenStudy (anonymous):

So what would it be for the 19 years?

OpenStudy (dumbcow):

saifoo how do you get 109,800 for slope? plug in the 14,19 into S(x)

OpenStudy (anonymous):

Come on..somebody please answer....

OpenStudy (saifoo.khan):

Oh wait.. my bad, lol.

OpenStudy (dumbcow):

@dumbunny S(x)=272,000 + 117,800x plug in 14 and 19 for x

OpenStudy (saifoo.khan):

S(x) = 272000+1450,000(14) 20572000 S(x) = 272000+1450,000(2010-1991) = 27822,000

OpenStudy (saifoo.khan):

here u go!

OpenStudy (anonymous):

Hey! I see the athlete salary question now! Seems I came in a little late :/

OpenStudy (saifoo.khan):

Oh hmm.

OpenStudy (anonymous):

dang, I wish I could get an answer to that athlete question.

OpenStudy (anonymous):

I see S(x) = 272,000 + 117,800....thats as far as I got.

OpenStudy (dumbcow):

haha c'mon you can figure it out. you know that for 2005, x = 14 and for 2010, x = 19 now where do you put the x in the equation? after the slope, right? the slope is 117,800 so multiply x(insert number here) by the slope then add 272,000

OpenStudy (anonymous):

No, I can't and I can't move to the next question until I get this one right. So I am pretty much screwed.